OpenStudy (anonymous):

A hunting lodge has enough fuel to keep 20 rooms heated for fourteen days. If the lodge decides to save fuel by turning off the heat in 5 unoccupied rooms, and each room requires the same amount of fuel to heat it, how many extra FULL days will the fuel supply last?

OpenStudy (anonymous):

I guess you could do a proportion: 20/14 = 15/x 20x = 210 /20 on both sides x = 10.5 So they will last 10 extra full days I believe

OpenStudy (anonymous):

just kidding i messed up

OpenStudy (anonymous):

The fuel is enough for 20*14 = 280 room-days. If you're heating 15 rooms, you can sustain heat for 280/15 = 18.6666... days. That's four extra FULL days.

OpenStudy (anonymous):

Does this make sense?

OpenStudy (anonymous):

@inkyvoyd @yomamabf Medal?